OpenStudy (anonymous):

an 80000kg airliner is flying at 900 kilometers/hour at a height of 10 kilometers. what is the approximate total energy?

OpenStudy (anonymous):

Total Energy = Potential Energy (PE) + Kinetic Energy(KE) Potential Energy is defined as the amount of energy that is stored at a certain height, equation: \[PE = mgh\] and Kinetic Energy is defined as the amount of energy that is in motion, equation: \[KE = \frac{ 1 }{ 2 }mv ^{2}\] where "m" is the mass, "v" is your velocity, "g" is your gravitational constant, and "h" as the height. You have all the variables given to you in the problem now just plug and chug. \[TotalEnergy = PE + KE\]

OpenStudy (anonymous):

Also, remember to switch the units of your variables from "kilometers" to "meters" before doing these equations, or you will end up with wrong answer.

OpenStudy (anonymous):

PE=(80000)(9.8)(10000m)=7840000000 KE=(1/2)(80000)(250m/s^2)=1/2(80000)(62500)=2500000000 TotalPower = 1.034 x 10^10? Im sure im not doing something right

OpenStudy (anonymous):

The answers are in MJ ie a)250MJ b)500MJ c)1,000MJ d)5,000MJ e)10,000MJ

OpenStudy (anonymous):

The key word here is approximate. Your answer is correct. But remember the unit for MJ is pretty much 1 joule x 1,000,000 or 1 Mega joule. Think about it closely. What does 1.034 x 10^10 really represent?

OpenStudy (anonymous):

I cant figure out how much 1 joule is

OpenStudy (anonymous):

A joule is an amount of energy, which the definition is not useful as of right now. The main point is remember about your scientific notations. You have the right answer, but you are assuming it's not right because it's not what you "exactly" see in the answers. But once again, they key word here is approximate, in which we choose an answer that is near ours. Remember "1 MJ" = 1,000,000 joules

OpenStudy (anonymous):

No, 10,000 + 1,000,000 = 1,010,000.

OpenStudy (anonymous):

oops i meant 10,000*1,000,000 or 1x10^10 / 1,000,000 is 10,000

OpenStudy (anonymous):

Yes. Good job. Now look at your answers and the answer you received. And choose the right one =)

OpenStudy (anonymous):

Thank you for the help I really appreciate it and I understand how to do the problem.

OpenStudy (anonymous):

You're very welcome. Cheers.